OpenStudy (anonymous):

HELP PLEASE HELP! ASAP. ∠x and ∠y are complementary. If m∠x = a + 3 and m∠y = 2a + 6, what is the measure of ∠x?

OpenStudy (anonymous):

do you know what "complementary" means?

OpenStudy (anonymous):

It adds up to 90

OpenStudy (anonymous):

ok good, so lets add them and set the result equal to 90

OpenStudy (anonymous):

\[\large a+3+2a+6=90\]

OpenStudy (anonymous):

know how to solve this one?

OpenStudy (anonymous):

combine like terms on the left and get \[3a+9=90\] then subtract 9 and get \[3a=81\] divide by 3 and get \(a=27\) but don't forget we solved for \(a\) and you still need \(x\)

OpenStudy (anonymous):

since \(x=a+3\) and \(a=27\) you have \(x=27+3=30\)

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

confused?

OpenStudy (anonymous):

Totes. What do you do after that?!

OpenStudy (anonymous):

ok, you got that they had to add up to 90 right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so when you add them you get \[a+3+2a+6=90\] ok so far?

OpenStudy (anonymous):

Yes, after that?

OpenStudy (anonymous):

combine like terms \(a+2a=3a\) and \(3+6=9\) so \[a+3+2a+6=3a+9\]

OpenStudy (anonymous):

this gives you \[3a+9=90\]

OpenStudy (anonymous):

ok to that step?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok good, now we have to solve for \(a\)

OpenStudy (anonymous):

\[3a+9=90\] subtract \(9\) from both sides to get \[3a=81\] then divide both sides by \(3\) to find \[a=27\] ok so far?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

now don't forget the question says "what is the measure of angle \(x\)?"

OpenStudy (anonymous):

we know that \(a=27\) and we also know that \(x=a+3\) replace \(a\) by \(27\) and get \[x=27+3\]

OpenStudy (anonymous):

I think I got it, is it 30?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw, hope steps are clear

OpenStudy (anonymous):

They sure were!